# Projecting Assumptions: The Duality Between Sparse Autoencoders and Concept Geometry

### v1

@@ E.1 Separability Experiment
Figure 9: Evolution of normalized MSE with training iterations for various SAEs on the separability experiment. Color intensity is proportional to $L_{0}$ (darker colors imply more dense SAE latents).
@@
Figure 10: Normalized MSE (normalized with variance of data), Stable ranks (of data correlations, latent correlations matrices), and fraction of dead latents as a function of sparsity ($L_{0}$) for the separability experiment (Sec. 5.1)
@@
Fig. 9 shows the evolution of normalized MSE (NMSE- MSE normalized by the variance of data) with training iterations for each SAE, for different levels of sparsity. Note that denser representations (higher $L_{0}$ and thus darker colors in Fig. 9) lead to lower NMSE. While all SAEs end up at similar levels of NMSE, their ability to extract concepts from data is markedly different (as described in Sec. 4). A per-concept breakdown of training dynamics is shown in Fig. 11. For comparison, this figure also includes the mean of the squared norm of each concept (which equals MSE if the SAE predicts the origin for all inputs), variance of each concept (which equals MSE if the SAE predicts the mean of each concept). Thus, SAEs whose MSE saturates at the concept variance are likely to be predicting the mean of the concept for all points, whereas when MSE goes below concept variance, the SAE explains within-concept variance. Also shown in gray is MSE with respect to the center of each concept, which ideally must match concept variance if the SAE reconstructs all points (which is observed in most cases).
@@
In Fig. 10, final NMSE as a function of sparsity ($L_{0}$) shows that while all SAEs have comparable MSE-sparsity curves at dense representations (high $L_{0}$), TopK’s NMSE goes down significantly more than others. This is a consequence of TopK learning a redundant solution, by just using two latents as an orthogonal basis to represent all data. Fraction of dead latents show large numbers of dead latents at high sparsity levels for ReLU, JumpReLU and TopK, with this going down (exponentially) as representations become more dense. However, SpaDE shows significantly fewer dead latents at all levels of sparsity. Stable ranks of cosine similarities between latent representations of pairs of data points (data corr.), and between pairs of latents across all data points (latent corr.) show that SpaDE has very high stable ranks, indicating high specialization of latents. The other SAEs have comparable stable ranks, all much lower than the desirable stable rank of $6$ (equal to the number of clusters in data).
@@
The SAE latent activation profiles for each concept are shown as histograms in Fig. 12. While variations exist across concepts, there is a common structure to the profiles for each SAE (SpaDE appears pointy, indicating a second mode other than zero).
@@
Cosine similarities between latent representations of pairs of data points are shown for different levels of sparsity in Fig. 13. Notice that SpaDE has the lowest cross-concept correlations of all SAEs, and these correlations do not decrease much especially in ReLU and JumpReLU. The corresponding figure with similarities between pairs of latents across all datapoints is in Fig. 14. Here, the number of dead latents increases with increasing sparsity, leading to very few active latents (only active latents are shown). Broadly, note the decrease in co-occurrences with increase in sparsity- also note how ReLU and JumpReLU result in newer correlation structures with greater sparsity.
@@
Figure 11: Training dynamics for each concept (column) across SAEs (rows) for separability experiment: colored solid lines are MSE, with intensity of color proportional to $L_{0}$. Gray lines show MSE of SAE predictions with respect to the center of each cluster; intensity is again proportional to $L_{0}$. . Black dotted line shows the mean squared norm of each cluster, which would equal the MSE if the SAE predicted the origin for all datapoints. Red dotted line shows variance of each cluster, which again equals MSE if an SAE predicts the center of the cluster. Note that when a model reconstructs data well, MSE wrt cluster center equals the variance of the cluster (as observed here)
@@
Figure 12: Histogram of latent representations for each concept of various SAEs on the separability experiment.
@@
Figure 13: Data correlations for various sparsity levels on the separability experiment: Pairwise cosine similarities between SAE latent representations of datapoints. White lines separate different concepts.
@@
Figure 14: Latent correlations for various sparsity levels on the separability experiment: Pairwise cosine similarities: pairwise cosine similarities between different SAE latents, computed across data from all concepts.
@@ E.2 Heterogeneity Experiment
The overall training dynamics (on data from all concepts) is shown in Fig. 16- note, again, that for low sparsity (high $L_{0}$, darker color) all SAEs reach similar levels of NMSE, but differ for higher sparsity levels. The per-concept breakdown of MSE, and comparison with mean squared norm, concept variance and MSE with respect to the center of each concept is in Fig. 11 . The kink in gray lines is precisely the point where the model transitions from learning to represent the mean, to learning to explain the within-concept variance, clearly demonstrating two phases in learning: learning the right scale for the data (since initial model predictions may not match the true scale of data), thereby predicting the mean well, followed by learning the distribution of the data.
@@
Fig. 18 shows latent activation profiles for each concept and each SAE ($k=32$ in TopK). Since TopK with $k=32$ cannot allocate enough latents for large intrinsic dimension concepts, it increases activations on smaller number of concepts instead. Cosine similarities between SAE latent representations for pairs of data points, and pairs of latents across all datapoints, is shown for varying levels of sparsity in Fig. 19, 20 respectively. All SAEs (except JumpReLU) do a decent job at reducing correlations between pairs of data points, but in the latent correlation plots, we see how TopK fails to adaptively allocate latents to heterogenous concepts, especially at moderate levels of sparsity, while the other SAEs do well- have different sized blocks in block-structured matrix.
@@
Figure 15: Normalized MSE (normalized with variance of data), Stable ranks (of data correlations, latent correlations matrices), and fraction of dead latents as a function of sparsity ($L_{0}$) for the heterogeneity experiment (Sec. 5.2)
@@
Figure 16: Evolution of normalized MSE with training iterations for various SAEs on the heterogeneity experiment. Color intensity is proportional to $L_{0}$ (darker colors imply more dense SAE latents).
@@
Figure 17: Training dynamics for each concept (column) across SAEs (rows) for heterogeneity experiment: colored solid lines are MSE, with intensity of color proportional to $L_{0}$. Gray lines show MSE of SAE predictions with respect to the center of each cluster; intensity is again proportional to $L_{0}$. . Black dotted line shows the mean squared norm of each cluster, which would equal the MSE if the SAE predicted the origin for all datapoints. Red dotted line shows variance of each cluster, which again equals MSE if an SAE predicts the center of the cluster. Note that when a model reconstructs data well, MSE wrt cluster center equals the variance of the cluster (as observed here)
@@
Figure 18: Histogram of latent representations for each concept of various SAEs on the heterogeneity experiment.
@@
Figure 19: Data correlations for various sparsity levels on the heterogeneity experiment: Pairwise cosine similarities between SAE latent representations of datapoints. White lines separate different concepts.
@@
Figure 20: Latent correlations for various sparsity levels on the heterogeneity experiment: Pairwise cosine similarities: pairwise cosine similarities between different SAE latents, computed across data from all concepts.
@@ E.3 Formal Language Experiments
In this section, we report several more results in the formal language experimental setup. Specifically, we show how with changing sparsity of the latent code, fidelity metrics, e.g., normalized MSE scales changes and stable rank of both data and latent correlations changes (Fig. 21); how monosemanticity changes, i.e., how F1 scores averaged across latents and the concept they achieve maximum F1 score on change (indicating their specialization to that concept) (Fig. 22 Left); and how percentage of dead latents change (Fig. 22 Right). These results are repeated at the concept-level, i.e., at the level of parts-of-speech, in Figs. 2324. Inline with results on heatmaps demonstrating correlation between sparse codes of samples from different concepts and between vector denoting which samples a given latent activates for, we retrieve results in Fig. 2526. The results above are perfectly inline with our findings from the main paper, e.g., that SpaDE achieves highly monosemantic features. The new and intriguing results involve demonstrations of how effective SpaDE can be at discerning position of a concept (part-of-speech) in a sentence, when compared to other protocols which learn a more uniform representation.
@@
Figure 21: Normalized MSE and Stable ranks as a function of sparsity in the Formal Language setup.
Figure 22: Monosemanticity (F1 scores averaged over latents) and fraction of dead latents as a function of sparsity for different SAEs in the Formal Language setup.
Figure 23: Normalized MSE decomposed by concepts (parts-of-speech) and plotted as a function of sparsity in the Formal Language setup.
Figure 24: Percentage of Dead Latents decomposed by concepts (parts-of-speech) and plotted as a function of sparsity in the Formal Language setup. Note that in such a concept-conditioned count of dead latents, one ends up counting both the latents that are always inactive and ones that are inactive for the specific concept under consideration.
Figure 25: Correlation between sparse codes of different concepts (parts-of-speech) in the Formal Language setup. Datapoints for different concepts are sorted according to which concept they come from (using a predefined order on the parts-of-speech) and according to their position in a sentence, hence highlighting position dependence. Lines demarcate boundaries at which tokens corresponding to different concepts start / end.
Figure 26: Correlation between which datapoints a latent activates for in the Formal Language setup. Latents are sorted according to which concept (part-of-speech) they most strongly activated for (as measured using F1-score). White lines demarcate boundaries at which latents of different concepts start / end.
Figure 27: 2D PCA visualization of sparse codes corresponding to different concepts (parts-of-speech).
Figure 28: 3D PCA visualization of sparse codes corresponding to different concepts (parts-of-speech).
Figure 29: 2D PCA visualization of a matrix whose elements capture which tokens a latent activates for. That is, which concepts (parts-of-speech) the latent is specialized towards, if any.
Figure 30: 3D PCA visualization of a matrix whose elements capture which tokens a latent activates for. That is, which concepts (parts-of-speech) the latent is specialized towards, if any.
@@ E.4 Vision Experiment
In this section, we show, visually, the concepts SpaDE has learnt in the vision experiment, by visualizing feature attribution maps for inputs from each class from Imagenette. We perform this visualization for the top concepts for each class for five classes- Tench (Fig. 31), Chainsaw (Fig. 32), Church (Fig. 33), Golf (Fig. 34) and Springer (Fig. 35)).
@@
Figure 31: Feature Attribution maps for monosemantic latents from SpaDE on the Tench class
@@
Figure 32: Feature Attribution maps for monosemantic latents from SpaDE on the Chainsaw class
@@
Figure 33: Feature Attribution maps for monosemantic latents from SpaDE on the Church class
@@
Figure 34: Feature Attribution maps for monosemantic latents from SpaDE on the Golf class

### v2

@@ E.1 Separability Experiment
Figure E.1: Evolution of normalized MSE with training iterations for various SAEs on the separability experiment. Color intensity is proportional to $L_{0}$ (darker colors imply more dense SAE latents).
@@
Figure E.2: Normalized MSE (normalized with variance of data), Stable ranks (of data correlations, latent correlations matrices), and fraction of dead latents as a function of sparsity ($L_{0}$) for the separability experiment (Sec. 5.1)
@@
Fig. E.1 shows the evolution of normalized MSE (NMSE- MSE normalized by the variance of data) with training iterations for each SAE, for different levels of sparsity. Note that denser representations (higher $L_{0}$ and thus darker colors in Fig. E.1) lead to lower NMSE. While all SAEs end up at similar levels of NMSE, their ability to extract concepts from data is markedly different (as described in Sec. 4). A per-concept breakdown of training dynamics is shown in Fig. E.3. For comparison, this figure also includes the mean of the squared norm of each concept (which equals MSE if the SAE predicts the origin for all inputs), variance of each concept (which equals MSE if the SAE predicts the mean of each concept). Thus, SAEs whose MSE saturates at the concept variance are likely to be predicting the mean of the concept for all points, whereas when MSE goes below concept variance, the SAE explains within-concept variance. Also shown in gray is MSE with respect to the center of each concept, which ideally must match concept variance if the SAE reconstructs all points (which is observed in most cases).
@@
In Fig. E.2, final NMSE as a function of sparsity ($L_{0}$) shows that while all SAEs have comparable MSE-sparsity curves at dense representations (high $L_{0}$), TopK’s NMSE goes down significantly more than others. This is a consequence of TopK learning a redundant solution, by just using two latents as an orthogonal basis to represent all data. Fraction of dead latents show large numbers of dead latents at high sparsity levels for ReLU, JumpReLU and TopK, with this going down (exponentially) as representations become more dense. However, SpaDE shows significantly fewer dead latents at all levels of sparsity. Stable ranks of cosine similarities between latent representations of pairs of data points (data corr.), and between pairs of latents across all data points (latent corr.) show that SpaDE has very high stable ranks, indicating high specialization of latents. The other SAEs have comparable stable ranks, all much lower than the desirable stable rank of $6$ (equal to the number of clusters in data).
@@
The SAE latent activation profiles for each concept are shown as histograms in Fig. E.4. While variations exist across concepts, there is a common structure to the profiles for each SAE (SpaDE appears pointy, indicating a second mode other than zero).
@@
Cosine similarities between latent representations of pairs of data points are shown for different levels of sparsity in Fig. E.5. Notice that SpaDE has the lowest cross-concept correlations of all SAEs, and these correlations do not decrease much especially in ReLU and JumpReLU. The corresponding figure with similarities between pairs of latents across all datapoints is in Fig. E.6. Here, the number of dead latents increases with increasing sparsity, leading to very few active latents (only active latents are shown). Broadly, note the decrease in co-occurrences with increase in sparsity- also note how ReLU and JumpReLU result in newer correlation structures with greater sparsity.
@@
Figure E.3: Training dynamics for each concept (column) across SAEs (rows) for separability experiment: colored solid lines are MSE, with intensity of color proportional to $L_{0}$. Gray lines show MSE of SAE predictions with respect to the center of each cluster; intensity is again proportional to $L_{0}$. . Black dotted line shows the mean squared norm of each cluster, which would equal the MSE if the SAE predicted the origin for all datapoints. Red dotted line shows variance of each cluster, which again equals MSE if an SAE predicts the center of the cluster. Note that when a model reconstructs data well, MSE wrt cluster center equals the variance of the cluster (as observed here)
@@
Figure E.4: Histogram of latent representations for each concept of various SAEs on the separability experiment.
@@
Figure E.5: Data correlations for various sparsity levels on the separability experiment: Pairwise cosine similarities between SAE latent representations of datapoints. White lines separate different concepts.
@@
Figure E.6: Latent correlations for various sparsity levels on the separability experiment: Pairwise cosine similarities: pairwise cosine similarities between different SAE latents, computed across data from all concepts.
@@ E.2 Heterogeneity Experiment
The overall training dynamics (on data from all concepts) is shown in Fig. E.8- note, again, that for low sparsity (high $L_{0}$, darker color) all SAEs reach similar levels of NMSE, but differ for higher sparsity levels. The per-concept breakdown of MSE, and comparison with mean squared norm, concept variance and MSE with respect to the center of each concept is in Fig. E.3 . The kink in gray lines is precisely the point where the model transitions from learning to represent the mean, to learning to explain the within-concept variance, clearly demonstrating two phases in learning: learning the right scale for the data (since initial model predictions may not match the true scale of data), thereby predicting the mean well, followed by learning the distribution of the data.
@@
Fig. E.10 shows latent activation profiles for each concept and each SAE ($k=32$ in TopK). Since TopK with $k=32$ cannot allocate enough latents for large intrinsic dimension concepts, it increases activations on smaller number of concepts instead. Cosine similarities between SAE latent representations for pairs of data points, and pairs of latents across all datapoints, is shown for varying levels of sparsity in Fig. E.11, E.12 respectively. All SAEs (except JumpReLU) do a decent job at reducing correlations between pairs of data points, but in the latent correlation plots, we see how TopK fails to adaptively allocate latents to heterogenous concepts, especially at moderate levels of sparsity, while the other SAEs do well- have different sized blocks in block-structured matrix.
@@
Figure E.7: Normalized MSE (normalized with variance of data), Stable ranks (of data correlations, latent correlations matrices), and fraction of dead latents as a function of sparsity ($L_{0}$) for the heterogeneity experiment (Sec. 5.2)
@@
Figure E.8: Evolution of normalized MSE with training iterations for various SAEs on the heterogeneity experiment. Color intensity is proportional to $L_{0}$ (darker colors imply more dense SAE latents).
@@
Figure E.9: Training dynamics for each concept (column) across SAEs (rows) for heterogeneity experiment: colored solid lines are MSE, with intensity of color proportional to $L_{0}$. Gray lines show MSE of SAE predictions with respect to the center of each cluster; intensity is again proportional to $L_{0}$. . Black dotted line shows the mean squared norm of each cluster, which would equal the MSE if the SAE predicted the origin for all datapoints. Red dotted line shows variance of each cluster, which again equals MSE if an SAE predicts the center of the cluster. Note that when a model reconstructs data well, MSE wrt cluster center equals the variance of the cluster (as observed here)
@@
Figure E.10: Histogram of latent representations for each concept of various SAEs on the heterogeneity experiment.
@@
Figure E.11: Data correlations for various sparsity levels on the heterogeneity experiment: Pairwise cosine similarities between SAE latent representations of datapoints. White lines separate different concepts.
@@
Figure E.12: Latent correlations for various sparsity levels on the heterogeneity experiment: Pairwise cosine similarities: pairwise cosine similarities between different SAE latents, computed across data from all concepts.
@@ E.3 Formal Language Experiments
In this section, we report several more results in the formal language experimental setup. Specifically, we show how with changing sparsity of the latent code, fidelity metrics, e.g., normalized MSE scales changes and stable rank of both data and latent correlations changes (Fig. E.13); how monosemanticity changes, i.e., how F1 scores averaged across latents and the concept they achieve maximum F1 score on change (indicating their specialization to that concept) (Fig. E.14 Left); and how percentage of dead latents change (Fig. E.14 Right). These results are repeated at the concept-level, i.e., at the level of parts-of-speech, in Figs. E.15E.16. Inline with results on heatmaps demonstrating correlation between sparse codes of samples from different concepts and between vector denoting which samples a given latent activates for, we retrieve results in Fig. E.17E.18. The results above are perfectly inline with our findings from the main paper, e.g., that SpaDE achieves highly monosemantic features. The new and intriguing results involve demonstrations of how effective SpaDE can be at discerning position of a concept (part-of-speech) in a sentence, when compared to other protocols which learn a more uniform representation.
@@
Figure E.13: Normalized MSE and Stable ranks as a function of sparsity in the Formal Language setup.
Figure E.14: Monosemanticity (F1 scores averaged over latents) and fraction of dead latents as a function of sparsity for different SAEs in the Formal Language setup.
Figure E.15: Normalized MSE decomposed by concepts (parts-of-speech) and plotted as a function of sparsity in the Formal Language setup.
Figure E.16: Percentage of Dead Latents decomposed by concepts (parts-of-speech) and plotted as a function of sparsity in the Formal Language setup. Note that in such a concept-conditioned count of dead latents, one ends up counting both the latents that are always inactive and ones that are inactive for the specific concept under consideration.
Figure E.17: Correlation between sparse codes of different concepts (parts-of-speech) in the Formal Language setup. Datapoints for different concepts are sorted according to which concept they come from (using a predefined order on the parts-of-speech) and according to their position in a sentence, hence highlighting position dependence. Lines demarcate boundaries at which tokens corresponding to different concepts start / end.
Figure E.18: Correlation between which datapoints a latent activates for in the Formal Language setup. Latents are sorted according to which concept (part-of-speech) they most strongly activated for (as measured using F1-score). White lines demarcate boundaries at which latents of different concepts start / end.
Figure E.19: 2D PCA visualization of sparse codes corresponding to different concepts (parts-of-speech).
Figure E.20: 3D PCA visualization of sparse codes corresponding to different concepts (parts-of-speech).
Figure E.21: 2D PCA visualization of a matrix whose elements capture which tokens a latent activates for. That is, which concepts (parts-of-speech) the latent is specialized towards, if any.
Figure E.22: 3D PCA visualization of a matrix whose elements capture which tokens a latent activates for. That is, which concepts (parts-of-speech) the latent is specialized towards, if any.
@@ E.4 Vision Experiment
In this section, we show, visually, the concepts SpaDE has learnt in the vision experiment, by visualizing feature attribution maps for inputs from each class from Imagenette. We perform this visualization for the top concepts for each class for five classes- Tench (Fig. E.23), Chainsaw (Fig. E.24), Church (Fig. E.25), Golf (Fig. E.26) and Springer (Fig. E.27)).
@@
Figure E.23: Feature Attribution maps for monosemantic latents from SpaDE on the Tench class
@@
Figure E.24: Feature Attribution maps for monosemantic latents from SpaDE on the Chainsaw class
@@
Figure E.25: Feature Attribution maps for monosemantic latents from SpaDE on the Church class
@@
Figure E.26: Feature Attribution maps for monosemantic latents from SpaDE on the Golf class